the students on the plaid team sold cookies to raise money. Peanut Butters cookies sold best. The students sold half as many chocolate chip cookies as peanut butter cookies. Twenty dozen cookies were sold altogether. How many cookies of each type were sold?
Answer by lwsshak3(11628)

the students on the plaid team sold cookies to raise money. Peanut Butters cookies sold best. The students sold half as many chocolate chip cookies as peanut butter cookies. Twenty dozen cookies were sold altogether. How many cookies of each type were sold?
***
12*20=240=total number of cookies sold
let x=number of chocolate chip cookies sold
240-x=number of chocolate peanut butter cookies sold
..
240-x=2x
3x=240
x=80
240-x=160
number of chocolate chip cookies sold=80
number of chocolate peanut butter cookies sold=160